2022 August/September Restart OUTLINE

UT September 2-11, 13, 21-22 2022 - Engineering On-Sky Restart Block

Instruments Released for Science

  • PEPSI-PFU - as of 9-3-2022
  • PEPSI-Pol - as of 9-5-2022
  • LBCs - as of 9-6-2002
    • a few tasks still outstanding (superfocus sequence) - but otherwise all else performs OK
  • LUCIs
  • MODS

Breakdown of Dates

  • Night given in UT, Daytime tasks in local time
  • UT Sep 02-07 (Sep 01-06 local time) - Full Nights
  • UT Sep 08 (Sep 07 local time) - Split Night w/PETS
    • Handover to PEPSI for reconfiguration UT 04:15 (21:15 local time)
    • Transit Start UT 04:45 (21:45 local time)
    • Transit End UT 08:15 (01:15 local time)
    • Handover from PEPSI to LBTO Eng-Restart (Reconfig) UT 08:16 (01:16 local time)
  • UT Sep 09-11 (Sep 08-10 local time) - Full Nights
  • UT Sep 13 (Sep 12 local time) - Split Night w/PETS
    • Handover to PEPSI for reconfiguration UT 03:50 (20:50 local time)
    • Transit Start UT 04:20 (09:20 local time)
    • Transit End UT 08:21 (01:21 local time)
    • Handover from PEPSI to LBTO Eng-Restart (Reconfig) UT 08:22 (01:22 local time)
  • UT Sep 21-22 (Sep 20-21) - Full Nights

General Scope of Activities

  • Closed Dome/Daytime Instrument Checks (each instrument must be released before night-time operations)
  • Telescope Wake-Up
  • Telescope Infrastructure
    • AGW1/2
    • M3 SX/DX
  • Checkout of Facility Instruments
    • Exported Scripts and OT files located in /home/LBTO/Service_Observing/Restart_2022 on OBS-N (64bit and obs-32bit) - by instrument name
    • LBC
    • MODS
    • LUCIs
    • PEPSI
      • PFU checkout
      • Pol checkout - high priority due to Pol run starting UT Sep 12
  • LBTI Restart (1-2 hr)
    • Close loop w/WFS only during first 10 days of restart (science cameras may not be cold enough)
    • Close loop w/science camera after UT Sep 11
  • AO (SX) Restart (2 x 0.5 nights)
    • Daytime calibraitons required

Key Dates & Information (Local Time)

  • DONE SX M3 Install 8/21
  • DONE DX M3 Already on telescope
  • DONE AGW-1 Install 8/23
  • DONE AGW-2 Already on telescope
  • DONE LUCI-1 Install 8/25
    • DONE cool-down 8/29 evening
  • DONE LUCI-2 already on telescope
    • DONE cool-down 8/29 evening
  • LBTI activities complete by ?
    • cold by ?
  • DONE LBC-B Install 8/31
    • TMS install 9/6-9/8
    • Software checks/updates through 9/9
  • DONE LBC-R Install 8/31
    • TMS install 9/6-9/8
    • Software checks/updates through 9/8
  • DONE SX AdSec Install - 8/24
  • DONE PEPSI SX Polarimater Install - 8/29
  • DONE PEPSI DX Polarimeter Install - 8/29
  • DONE DX AdSec Install - 8/31
  • DONE PESPI PFU Intervention (AIP team) - 8/29
  • DONE MODS-1 install 9/20
  • DONE MODS-2 install 9/20

Closed Dome/Daytime Work

LBC (2 hrs)

  • DONE Filter Functional checks
  • DONE Rotator Functional checks
  • DONE Test Bias Exposures

TMS Testing for LBCs (4hrs)

  • DONE Requires moving telescope - 9/9-9/10

MODS (Sep 20)

  • DONE Functional Checkout (2hrs) - telescope at zenith
    • M1
    • M2
  • DONE Test Exposures (sieve and internal flat snaps) - telescope at zenith
    • M1
    • M2
  • DONE Internal Calibrations (2hrs) - telescope at zenith
    • Take 5" slit images to check the reference position for alignments: red-only, blue-only
    • Dual mode slit flats to check for debris
    • Red-only 5" slit flats
    • M1
      • Grating Mode
      • Prism Mode
      • Imaging Mode
    • M2
      • Grating Mode
      • Prism Mode
      • Imaging Mode
  • DONE MODS Mask Exchange - telescope at zenith
    • DONE M1
    • DONE M2
  • MODS Focus Checks - telescope at zenith
    • M1
    • M2
  • DONE AGw5/AGw6 aliveness check (1 hr)
    • take images with all cameras (could be closed-dome preset)
    • move stages
  • DONE Test scripts on obs1, obs2
    • modsWake
    • modsSleep
    • modsSimSnap
    • modsSieveSnap
  • Test operation of MODS on obs1, obs2 under all partner accounts.
    • DONE Run up the GUIs
      • odd message on OBS1 and OBS2 when closing MODS2 GUI
      • QThread: Destroyed while thread is still running
    • DONE modsDisp
    • DONE modsView
    • run modsSieveSnap (as a check on sending commands, taking data, displaying the images)
    • run a mock modsAlign
      • send offsets (MODS authorized, TCS enabled).
    • Run modsView, IDL quick reduce pipeline
  • Verify that updatepointing only works for detxy, not for radec coordinates

LUCIs (5-6 days)

  • Basic Checks needed for release for engineering (and AO) work - 1 day i n total - 8/29
    • DONE non-MOS Functional Checkout on telescope L1 (cold) - 0.5 day - internal - no restrictions on telescope
    • DONE non-MOS Functional Checkout on telescope L2 (cold) - 0.5 day- internal - no restrictions on telescope
    • DONE MOS Functional Checkout on telescope L1 (cold) - 0.5 day - internal - no restrictions on telescope
    • DONE MOS Functional Checkout on telescope L2 (cold) - 0.5 day - internal - no restrictions on telescope
    • DONE PFS alignment and Camera and Grating refocus L1 - 1hr - requires L1 authorized and M3 aligned
    • DONE PFS alignment and Camera and Grating refocus L2- 1hr - requires L2 authorized and M3 aligned
  • Grating Wavelength Calibration - 3-5hr take data, 1-2 days process data (cal unit in beam - stationary telescope)
    • DONE G210 L1
    • DONE G210 L2
    • DONE G200 L1
    • DONE G200 L2
    • G040 L1 - data collected
    • G040 L2 - data collected
  • Flexure Checks - 3 days? - Requires moving telescope (moving instrument rotator and elevation drive) - Starting 8/29 (Requires basic checks to be completed first)
    • DONE G210/N1.8 - L1/L2
    • DONEG210/N3.75 - L1/L2
    • DONEG200/N3.75 - L1/L2
    • DONEG200/N1.8 - L1/L2
    • G040/N30 - L1/L2
    • DONEN1.8 imaging - L1/L2
    • DONE N3.75 imaging - L1/L2
    • DONE N30 Imaging - L1
    • N30 Imaging L2
  • Mask Exchange (Aug 30, Sep 2) - 4hrs each - telescope locked at zentih
    • DONE L1
    • DONE L2
  • LUCI Software checks after LUCIs released for engineering nighttime on-sky use
    • DONE verify lucix can run the LUCIs - swap fibers to lucix and back to luci - G.Eychaner - to be done week of 9/5
    • Replace LUCI-1 WEBIO with one configured with correct IP address (IT 8083) - G.Eychaner - to be done week of 9/19
    • DONE Test GEIRS abort - bug in similator - verify if it occurs or not with detector - G.Eychaner
    • Test Double-Buffering and overheads - D. Thompson - done after functional/calibration checkouts

LBTI (2 days)

  • Internal Checkout (0.5 day)
  • AdSec check out (1.5 day)

AO (9 days)

  • SX Pupil Wobble Test - 0.5 day - earliest can be done 8/30
    • Requires AGW-1 install
    • May Requires LUCI-1 Install
    • May Requires LUCI-1 Functional Checkout Imaging
  • DONE SX ASM Calibration of Feed Forward Matrix - 8/25 - 1 day
  • DONE SX ASM Flattening w/4D Interferometer - 8/26-8/29 - 4 days
    • DONE Seeing Limited Flat - 1 day - 8/26
    • DONE Diffraction limited Flat - 2 days - 8/27-8/28
    • DONE Contingency - 1 day
  • SX LUCI-1 AO Reconstructors - 4 days 8/31-9/4
  • DONE SX AO - verify closed dome - 3 days - week of 9/5
    • DONE Bin 1,2,3,4 w/ and w/o NCPA
    • DONE ESM Bin 1,2,4
    • Argos cal-unit only
  • DX tasks
    • No DX daytime work needed - all previous calibrations should work after reinstall

PEPSI (1 day for each focal station)

  • Internal funcitonal checkout
    • DONE PFU
    • DONE Polarimeters - ´╗┐ready for on-sky restart
    • DONE Calibration

Potential Conflicts/Issues

  • MODS
    • Both MODS installed on Sep 20 daytime
      • after install need 4-5 hours of functional checkout before going on-sky
      • after install need to do mask exchange (5 masks each MODS)
  • LUCI / TMS
    • L2 mask exchange Sep 1, coincides with software and other testing - if these TMS tests don't impact telescope or need motion, then OK
    • TMS requires moving telescope - could be done on Sep 2
  • LUCI Checkout / AO calibration
    • LUCI checkouts begin Aug 29 - SX AO Reconstructors may overlap with LUCI-1 exchange
    • DX ASM slated for install 8/31, if delayed, could impact L2 exchange on 9/1
    • No conflict with 4D and LUCI focal station
  • LBCs - functional checks could occur after installation, later LBCs installed more chance of interferring with LUCI (8/29 onwards) and AO (SX ASM install 8/26 testing after)
  • TMS/AO
    • DX ASM install on 8/31 - can this be done in parallel with TMS install on SX? No, zenith vs horizon conflict.

List of On-Sky Tasks

Telescope Wake-up (1 hr)

  • DONE Closed Dome Test Presets (multiple instruments)
    • DONE LBC
    • DONE LUCIs
    • DONE PEPSI
    • DONE MODS
  • On-sky presets (Track, Acquire, Active)
    • DONE LBC Track
    • DONE LUCIs Active, Acquire, Track
    • DONE PEPSI, PEPSI Pol - Active, Acquire
    • DONE MODS Active, Acquire, Track- Sep 20

Telecope Infrastructure

  • UPDATE - IDL AGW tasks on OBS-1 & OBS-2 - use "idl_agw" to launch an IDL window and run takss from there
  • AGW1/2 Alignment (LUCI-1) (1 night) Overview of AGW procedure
    • Files needed:
      • ssh -X into TO1 as telescope; then ssh - X as tcs@tcs1:
      • /home/telescope/TCS/Configuration/GCS/LUCI_L.cfg
      • image files: /lbt/data/share/tcs/lbto.YYYMMDD
      • oac@oac: /lbt/oacontrol/current/etc/oacontrol.conf
    • DONE Pupil alignment check - find a star and center it - need Bent Gregorian (1 hr) - need to do before AGW checkout
    • DONE verify SH spots illuminated (0.5 hr)
    • DONE align guiding hotspot to rotator center (1 hr)
    • DONE SX M3 Alignment LFBG - Critical (3 hrs) - requires a LUCI AGw - both sides can be done in parallel
      • Pupil wobble (2 hrs) - requires LUCI-1/AGW-1 to guide, collimate, SH spots well illuminated, and hotspot centered - can be done before AGW Rot ZP and Transform
    • DONE DX M3 Alignment RFBG Critical (3 hrs) - requires a LUCI AGw - both sides can be done in parallel
      • Pupil wobble (2 hrs) - requires LUCI-1/AGW-1 to guide, collimate, SH spots well illuminated, and hotspot centered - can be done before AGW Rot ZP and Transform
  • DONE SX M3 Alignment PFU - check
  • DONE DX M3 Alignment PFU - check
  • DONE Field Aberration (2 hrs) - High Priority (1hour) (SX more important than DX due to SX removal)
    • DONE SX w/LUCI
    • DONE DX w/LUCI
      • Just need both AGW1 and AGW2 on the telescope and checked to execute
      • requires AGW-1/AGW-2 or AGW5/6 to guide, collimate, SH spots well illuminated, hotspot centered - can be done before AGW Rot ZP and Transform
      • Requires < 1" to get good results
        • Don't hold out for 1", take in 1.2" to get first order results.
    • DONE SX data analyzed
    • DONE DX data analyzed
    • DONE SX correction installed
    • DONE DX correction installed
    • DONE SX w/LUCI check data collected and processed
    • DONE DX w/LUCI check data collected and processed

LBC (high)

  • DONE Collimate WRS/DOFPIA - (4 x 0.25 hrs)
  • Superfocus sequence -PA=0,180, 90, 270 - high (1 hr)
  • DONE Rotator Center Checks, start with PA=0, 180 (track mode), then 90,270 - high (1hr)
  • DONE Test Science OB - guiding - medium (1 hr)
    • DONE Sidereal
    • DONE Non-Sidereal OB
  • DONE Standard photometric fields (new ones ready) - requires clear skies (any seeing) - required for ZP calibrations
  • Spectro-photometric standards - requires clear skies (any seeing) - required for ZP calibrations
  • DONE Twilight Flats all filters - High (4 x 0.5 hrs - likely min two nights to complete)
    • The following are completed:
      • sloan-g LBC-B 0,180
      • sloan-i LBC-R 0,180
      • V-Bessell LBC-B 0,180
      • V-Bessell LBC-R 0,180
      • U-Bessell LBC-B 0,180
      • Y-Fan LBC-R 0,180
      • Uspec LBC-B 0,180
      • F972 LBC-R 0,180
      • B-Bessell LBC-B 0,180
      • R-Bessell LBC-R 0,180
      • r-sloan LBC-B 0,180
      • z-sloan LBC-R 0,180
      • sloan-r LBCB 0,180
      • sloan-r LBCR 0,180
      • I-Bessell LBCR 0,180
    • Missing:
      • I-Bessell LBCR 0,180
  • TMS with LBC (Heejoo, Olga and John H.) - 4hrs - earliest is UT 9/11 (local 9/10)
  • LBC Pointing Model - (3hr) - any conditions
    • TMS (active) - low priority
      • unsure if pointing model can be generated if run in active mode
      • timing problems - slewing vs TMS inputing corrections (JMH may have fixed this with his update to lbc_tms.py on 20210905), so we should at least try the active TMS pointing model
    • Standard method -medium priority
      • PtStarrs - please choose targets > 80 deg elevation (more than 1 or 2)
      • Run TMS passively - run 1 reference at the beginning
      • Need stable temperatures
      • Turn on PCS logging
      • Need to collimate at each star
      • Chip 2 only exposure
      • Analysis done by - O. Kuhn
      • File Location:
        • /home/lbto/Calib_OBs/PtStars_10m_OBs/VV
  • Cryostat tilt adjustment check
    • Use the superfoc data
    • Calculate shim thickness from data - off telescope
    • Iterations might be needed - 1hr per iteration

MODS (high) - UT Sep 21, Sep 22

  • DONE Collimate AGWs
  • DONE Check SH spots
  • DONE check MODS1/MODS2 guide probe wrt/rotator center
    • Files needed:
      • agw.ini located on mods1 and mods2 computers in: /home/mods/Config/MODS1 or /home/mods/Config/MODS2
    • M1 hotspot found to be 250,198, Hotspot is (x,y): 254.80, 201.00 - this is OK
    • M2 hotspot found to be 269, 217, Hotspot is (x,y): 267.00, 240.00 (there is a ~30pixel hardware limit on Y so this is OK)
  • DONE MODS Rotator Center measure
    • requires pointing mask (40" square field stop & dot-circle with radius 45") - high (0.5hr)
  • DONE MODS AGW 5/6 - Measure Rotator zero point (mask alignment) - high (0.5 hr) - mask exchange required
    • Target Masks:
      • ID 532732 - mods.4.g04d25ep.mms
      • M1 theta = 0.011
      • M2 theta = -0.034
      • No changes made to Rot ZPs for either MODS
    • ssh -X into TO1 as telescope; then ssh - X as tcs@tcs1
      • The files needed are
        • /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
  • MODS AGW Transform data - ~1" seeing - high (2hours to collect data on-sky, time during day to analyze, ~0.5 hr-1hr to verify on-sky) - must do RotZP First
    • Not done - RotZP unchanged - no indication need to check - can do if run out of other higher priority eng tasks
    • Lower Priority - skipped for now as RotZp are unchanges
    • Verify Tansform data
      • Update transform from measurements
        • stop AGW services (mod1 stop agw, mods2 stop agw)
        • new file copied to ~/Config/MODSN/ and renamed agw.ini
        • Restart the AGW services (mods1 start agw, mods2 start agw)
      • Collect transform data for verification
      • Measure Rot ZP with MODS MOS masks
      • Dither Test M52 or M71
      • Confirmation new transform is acceptable
  • Spectrophotometric standard all modes -High (to compare SX vs DX)
    • Grating
    • Prism
  • non-sidereal guiding
    • Imaging with dithers
    • Long-slit
  • Test Science OB
    • longslit
    • imaging
      • Test without syncoffset and large dither (+/-30")
        • got co-pointing warnings
      • Test with synoffset and large dither (+/-45")
  • Dithering Repeatability with Pointing Mask to check DET XY
  • Skyflats
  • Blind offset tests
  • Pointing Model - 2hr - low priority
  • GCS and IDL have opposite field correction (IT 8315) - Medium/High (1-1.5hrs)
    • Requires 1" or better
    • run IDL task field_collect (same as measuring field aberrations)
    • take LUCI data first to cross-check IDL vs GCS

LUCIs (high)

  • Rotator center
    • DONE L1
    • DONE L2
  • DONE rotator zeropoint (requires MOS mask) - (0.5 hr)
    • Masks:
      • ID 903033 - Stone O
      • ID 921934 - Stone A
      • iD 931027 - Stone D
    • To update results (if needed)
      • ssh -X into TO1 as telescope; then ssh - X as tcs@tcs1; The file location is in /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
    • Longslit Method (backup for no masks)
      • requires target field with two sources in slit, use alignment holes in longslit. Need LMS file with source in longslit (preferably two) - astrometric field ideal
  • DONE Collect transform data - (2hrs) - must do RotZP First

    • Target Fields:
      • Stone O 21 37 31 +00 31 55 Stone_O_m12.list
      • Stone P22 35 49 +01 31 33 Stone_P1.list
      • Stone A 00 54 02 -00 32 34 Stone_A3.list
    • DONE Process transform data
    • DONE Update transform
      • ´╗┐/lbt/oacontrol/current/etc/oacontrol.conf
      • rdwrconfig -u1 and rdwrconfig -u2
      • restart PCS and GCS
    • DONEverify transform data (1 hr)
      • data collection
      • 11/12 for LUCI1, 9/12 for LUCI-2
    • DONE Final AGW transform in place
    • DONE Rot ZP check
      • Results
        • LUCI-1
          • Stone O LS method
            • L1 0336-0338
            • seeing <1" on guiders
            • err 0.062", 0.172"
            • xshift = 0.217", yshift = 1.846", theta = -0.061
            • Repeated UT 09 06 22
            • Seeing 1.5" on guiders
            • err 0.061", 0.091"
            • xshift 0.221", yshift 0.003" theta = -0.057
          • Stone P LS method
            • L1 0026-0028
            • seeing < 1"
            • err 0.081, 0.0741
            • removed #37
            • xshift =0.375", yshift = -0.136", theta = 0.025
          • Stone A LS method
            • L1 0031-0033
            • seeing 1"
            • err 0.078", 0.080"
            • xshift = 0.369", yshift = -0.0407" theta = -0.003
        • LUCI-2
          • Stone O MOS mask
            • L2 0317-0322
            • Seeing <1" on guiders
            • eliminating #36, Err 0.068,0.138
            • shifts x=1.481", y=-3.490", theta =-0.134
            • removing #33 only changed delta theta=-0.12
          • Stone O LS method
            • L2 0315-0317
            • Seeing <1" on guiders
            • shifts x=1.32" , y=-2.94" , theta = -0.124
            • Repeated UT 09 06 22
            • seeing 1.5"-2" on guiders
            • err 0.083,0.089
            • xshift = 1.12" yshift = 1.19", theta= -0.124"
          • Stone A MOS mask
            • L2 0323-0326
            • Seeing 0.7" in guiders
            • all stars good Err 0.056, 0.057
            • Xshift = 1.737", Yshift = -3.462", theta=-0.079
          • Stone D MOS mask
            • L2 0415-0419
            • poor seeing (1.7" on K-band images)
            • err 0.096",0.136"
            • x-shift =0.94", yshift = -3.27", theta = -0.071
          • Stone P LS method
            • L2 0024-0026
            • seeing <1"
            • err 0.099", 0.0091"
            • removed #36
            • xshift = 0.149", yshift = -3.585", theta = -0.062
          • Stone A LS method
            • L2 0029-0031
            • seeing 1"
            • err 0.0859", 0.069"
            • xshift = 1.233" yshift -3.365", theta = -0.073
      • Stone O is an outlier for both LUCIs, using the other fields all are < 0.1" theta - so no update will be made to PCSInstrument.conf for the RotZP
  • DONE Imaging check
    • Check was done via taking bino-photometric standards - dither in-field
    • dither in field (0.5hr)
    • dither out of field (0.5hr)
      • L1
      • L2
  • DONE non-sidereal Imaging Check
    • dither in field (0.5hr)
    • dither out of field (0.5hr)
  • check longslit - high (1 hr)
    • DONE Science-like OB for G210
    • Science-like OB for G200
  • check longslit - NS
    • Only 1 grating test needed if sidereal above is done
  • Twilight flats - low priority
  • Pointing Model - High (2hrs)
    • PtStarrs ~10th mag 12-20 stars
    • Poor seeing/clouds OK
  • DONE Photometric Standards - any seeing, must be clear
    • DONE L1
    • DONE L2

Mixed Mode Tests (0.5 night)

  • MODS/LUCI Mixed Mode
  • LBC/LUCI
  • LBC/MODS

LBTI (up to 0.5 night)

  • Close the loop w/WFS (no science camera needed, just to make sure AO functioning) - 2hrs
    • can be done after ASMs released - and during first restart block
  • Closer the loop w/Science camera- 2hrs
    • requires cameras to be cooled and ready - likely after initial restart block
  • Full restart tasks can be executed in Early October (UT 10/4, UT 10/8, or UT 10/9)

AO (0.5 night)

  • System checkout
    • OSCO
    • Diffraction limited verification - use AO OT file - select a star, run through script
      • Bin 1 star (brighter than OSCO)
      • Bin 2 star
      • Bin 3 star
      • Bin 4 star
      • Bin 4 star faintest (WFSmag ~17.5-18)
      • Non-sideral (any brightness)
    • ESM verification (select a target from the ESM OT file, like M92, M52, etc)
      • Bin 1
      • Bin 2
      • Bin 4

PEPSI (0.5 night -high)

  • DONE Verify PFU AGW 7/8 alignment
  • DONE Verfiy POL AGW 3/4 Alignment
  • DONE Flux standard for PEPSI-Pol
  • DONE Flux standard for PFU (pending fiber intervention)
  • Pointing Model - 2hr Medium
    • Remove retro-reflecting dichroic (to give full FOV)
    • Take images with PFU (Active mode) -not a requirement, images come "free"
    • Need ~28 stars (10th mag or so)
      • PtStars (IRAF) will generate stars based on LST
      • minimized long azimuth slews
    • Analysis with TPOINT - D. Thompson
    • Can do in poor seeing/clouds

Rainy Night Tasks

-- BarryRothberg - 15 Jul 2021

Support Files

MMS files:

Field MMS RA Dec MaskID Manufacture
Gaia 00+25 mods.2.g00d25ep.mms 00 00 00.139 +25 00 17.090 531713 2021
Gaia 20+25 mods.3.g2025ep2.mms 20 00 00.910 +25 00 28.100 531956 2021
Gaia 04+25 mods.4.g04d25ep.mms 03 59 58.669 +25 00 40.050 532732 2021

LMS Files:

Field RA (J2000) Dec (J2000) MASK ID
StoneO (new) 21 37 31.791 00 31 55.010 903033
StoneP (new) 22 35 49.616 01 31 33.950 933322
StoneA (new) 00 54 02.050 -00 32 34.460 921934

LMS Files with unboxed stars 1" longslit mask:

Field RA (J2000) Dec (J2000) Finder
M048stnOZP +975011 21 37 31.632 00 31 46.820 Stone O 1" longslit
M056stnPZP +912345 22 35 49.689 01 31 27.970 Stone P 1" longslit
M053stnAZP +991823 00 54 03.544 -00 32 12.120 Stone A 1" longslit
Scripts, LMS files, andfinder charts are located in the linked google drive. The OT RotZP _v2.xml contains the OB's used to generated the necessary scripts. Presets are sent with the normal LS_1.00arcsec slit mask. Rather than a longslit alignment, users should execute a MOS alignment, uploading manually the corresponding lms file. The RTD MOS mask alignment assumes that rather than slit #1 and #6 used across the top as is typically done, #1 is the central alignment hole in the bundle up top and #6 is the alignment hole in the bundle at the base of the mask. Once aligned, unboxed stars are used to refine the mask alignment. There are 2 stars located at the extremes of the slit that can be used for alignment verification. Andy questions, contact Jenny.

-- BarryRothberg - 22 Aug 2022
Topic revision: r47 - 26 Oct 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback